if 15 miles are covered in 2 hours than 90 miles are covered in how many time
Answer by stanbon(75887)

if 15 miles are covered in 2 hours than 90 miles are covered in how many time
-----
Use a proportion::
time/distance = time/distance
------
t/90 = 2/15
------
t = 90(2/15)
------
t = 12 hrs
===================
Cheers,
Stan H.
==================